OpenStudy (anonymous):

Three balls are numbered 1, 2, 3 and placed in a bag. A ball is chosen, its number recorded, then the ball placed back in the bag. This process is repeated twice more. The resulting 3 recorded numbers are added. Bob walks in the room after the three numbers have been recorded and is told only that the sum of the numbers is 6. Bob announces, "Oh, then they must all have been 2's." What is the probability that Bob is correc

OpenStudy (anonymous):

The sample space has 27 triplets. Like 1 1 1, 1 1 2, 1 1 3 etc. Figure out how many of them add up to 6 call that n P = 1/n

OpenStudy (anonymous):

Or start with the middle digit = 1, 2, or 3 If 1, then only 2 1 3 and 3 1 2 work If 2, then only 3 2 1 and 1 2 3 work There's a pattern here!

OpenStudy (anonymous):

Forgot 2 2 2

OpenStudy (anonymous):

I believe there are 7 triplets with this property.

OpenStudy (anonymous):

let me see. but also thanks for the info

OpenStudy (anonymous):

Yea there are 7

OpenStudy (anonymous):

so p= 1/7

OpenStudy (anonymous):

@telliott99 did you see what I wrote?

OpenStudy (anonymous):

back

OpenStudy (anonymous):

Yes that sounds good.

OpenStudy (anonymous):

@telliott99 thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Solution: The only possible sums that give 6 are \[1+2+3\] and \[2+2+2.\] Of the \[3^3=27\] total ways to draw three balls in order, there are \[3\cdot 2\cdot 1=6\] ways to draw 1, 2, and 3 in some order. There is also 1 way to draw 2, 2, and 2. There are 7 total ways to draw three balls that sum to 6. Of these 7 ways, only one of those ways is the draw 2, 2, and then 2. Therefore Bob is only correct in 1 of those 7 cases. The probability that Bob is correct is \[\frac{1}{7}\].